Solve
x^2y"- 2xy' + 2y=x^2+sin(5Inx)
1
Expert's answer
2020-10-21T15:56:27-0400

"\displaystyle x^2y"- 2xy' + 2y=x^2+\\sin(\\ln(5x))\\\\\n\nx = e^t\\\\\n\ne^{2t} y" - 2e^{t}y' + 2y = e^{2t} + \sin(5t) \\


y = f(\ln(x)) \\


\frac{\mathrm{d}y}{\mathrm{d}t} = \frac{f'(\ln(x))}{x}\\


\frac{\mathrm{d}^2y}{\mathrm{d}t^2} = \frac{-f'(\ln(x))}{x^2} + \frac{f''(\ln(x))}{x^2} = \frac{f"(\\ln(x)) - f'(\\ln(x))}{x^2}\\\\\n\n\n\ne^{2t}\\left(\\frac{f"(\ln(x)) - f'(\ln(x))}{x^2}\right) - 2e^{t}\frac{f'(\ln(x))}{x} + 2f(\ln(x)) = e^{2t} + \sin(5t) \\


e^{2t}\left(\frac{f"(\\ln(x)) - f'(\\ln(x))}{e^{2t}}\\right) - 2e^{t}\\frac{f'(\\ln(x))}{e^t} + 2f(\\ln(x)) = e^{2t} + \\sin(5t)\\\\\n\n\nf"(\ln(x)) - f'(\ln(x)) - 2f'(\ln(x)) + 2f(\ln(x)) = e^{2t} + \sin(5t) \\


f"(\\ln(x)) - 3f'(\\ln(x)) + 2f(\\ln(x)) = e^{2t} + \\sin(5t) \\\\\n\n\n\nf"(t) - 3f'(t) + 2f(t) = e^{2t} + \sin(5t) \\

f(t) = f_p + f_c\\


f_c \, \textsf{is the complementary factor while}\\f_p \, \textsf{is the particular integral}\\

\textsf{The auxiliary equation is}\\

m^2 - 3m + 2 = 0\\

(m - 1)(m - 2) = 0\\

m = 1, 2\\


f"(t) - 3f'(t) + 2f(t) = e^{2t} + \\sin(5t)\\\\\n\n\n\\textsf{Solving by the method of variation of parameters}\\\\\n\nW(t) \\, \\textsf{is the Wronskian of the solution to the DE}\\\\\n\nW(t) = e^{-\\int -3 \\, \\mathrm{d}t} = e^{3t}\\\\\n\n\nV_1(t) = -\\int e^t \\frac{(e^{2t} + \\sin(5t))}{e^{3t}} \\,\\mathrm{d}t = -\\int (1 + e^{-2t}\\sin(5t))) \\,\\mathrm{d}t\\\\\n\n\n\\int e^{-2t}\\sin(5t) \\, \\mathrm{d}t = -\\frac{2e^{-2t}\\sin(5t)}{29} -\\frac{5e^{-2t}\\cos(5t)}{29}\\\\\n\n\nV_1(t) = -t + \\frac{2e^{-2t}\\sin(5t)}{29} + \\frac{5e^{-2t}\\cos(5t)}{29}\\\\\n\n\\begin{aligned}\nV_2(t) &= \\int e^{2t} \\frac{(e^{2t} + \\sin(5t))}{e^{3t}} \\,\\mathrm{d}t \\\\&= \\int e^t + e^{-t}\\sin(5t) \\,\\mathrm{d}t \\\\&= e^t - \\frac{5}{26}e^{-t} \\cos(5t) - \\frac{1}{26} e^{-t}\\sin(5t)\\\\\n\\end{aligned}\\\\\n\n\\begin{aligned}\nf_p &= \\left(-t + \\frac{2e^{-2t}\\sin(5t)}{29} + \\frac{5e^{-2t}\\cos(5t)}{29}\\right)e^{2t} \\\\&+ \\left(e^t - \\frac{5}{26}e^{-t} \\cos(5t) - \\frac{1}{26} e^{-t}\\sin(5t)\\right)e^t \\\\&= -te^{2t}+ \\frac{2\\sin(5t)}{29} + \\frac{5\\cos(5t)}{29} + e^{2t} - \\frac{5}{26} \\cos(5t) - \\frac{1}{26}\\sin(5t) \\\\&= e^{2t}(1 - t) - \\frac{15}{784}\\cos(5t) + \\frac{23}{754}\\sin(5t)\\\\\n\\end{aligned}\\\\\n\n\nf"(t) - 3f'(t) + 2f(t) = e^{2t} + \sin(5t)\\


\therefore f(t) = Ae^t + Be^{2t} + e^{2t}(1 - t) - \frac{15}{784}\cos(5t) + \frac{23}{754}\sin(5t)\\


\therefore y = f(x) = Ax + Bx^2 + x^2(1 - \ln(x)) - \frac{15}{784}\cos(5\ln(x)) + \frac{23}{754}\sin(5\ln(x))\\

\textsf{Is the solution to the ODE}"